OpenStudy (anonymous):

Use the substitution method to find the solution(s) to the system of equations. Show all work for full credit. y = –x^2 y = x

OpenStudy (foolaroundmath):

y = -x^2 y = x Substituting y =x in first equation x = -x^2 => x^2 + x = 0 => x(x+1) = 0 => x = 0 or x+1 = 0 => x = 0 or -1 => y = x i.e. y = 0 or -1 So the solutions are (x,y) = (0,0) or (-1,-1)

OpenStudy (anonymous):

i think the solutions are (0, -1) though im so confused